OpenStudy (anonymous):

g^2+2/3g+1/9=0 solve the equation

OpenStudy (anonymous):

this will factor into (g+1/3)(g+1/3) = 0 so setting the factors equal to 0 g+1/3=0 g=-1/3

OpenStudy (anonymous):

would positive 1/3 be included in that

OpenStudy (anonymous):

nope, because the other factor is also g+1/3 this function touches the x-axis but does not cross it

OpenStudy (anonymous):

ok thanks